alextrejo373:

3. (08.02 LC)The table below shows the size of nine families selected at random from two neighborhoods in a large city:Family Size (in number of people)Which neighborhood appears to have a bigger family size? Explain your answer using two or three sentences. Make sure to use facts to support your answer. (10 points) Neighborhood Q 6 3 1 2 3 2 5 1 2 Neighborhood S 5 2 4 3 5 3 2 4 5

Florisalreadytaken:

Neighborhood Q 6 ; 3 ; 1 ; 2 ; 3 ; 2 ; 5 ; 1 ; 2 (9 families in total) Neighborhood S 5 ; 2 ; 4 ; 3 ; 5 ; 3 ; 2 ; 4 ; 5 (9 families in total as well) Which neighborhood appears to have a bigger family size? well to determine that, we have to find the average people per family Neighborhood Q= \(\frac{6 + 3 + 1 + 2 + 3 + 2 + 5 + 1 + 2}{9} \Rightarrow \frac{25}{9} \) Neighborhood S=\( \frac{5 + 2 + 4 + 3 + 5 + 3 + 2 + 4 + 5 }{9}\Rightarrow \frac{33}{9}\) so, which of the neighborhoods has got the bigger family size?